Question 4 30 pts ial a 1520° V {8-82 Vab n 152 - 240° V 152-120° V {8-82 {8-82 Vca b in þvoc с iaFind the line currents and line-to-line voltages. Round answer to 2 decimal places, add Os as needed. Express magnitude as p

Kindly do this clearly.
Thanks.
Answer #1

4 Given circuit can be redrawn as lad va 15LOU 88-38 a 154-240V © 154-16V į8-J8A Aube Ice Since the load balanced ; we con Trla Ibl= 1.325 (45-1200 Ibe = 1.325 _-75A 120 128 Ib Icd = 1-325 245°+120° NAY Ich 1-325 2165*A 4 The line currents are ial =4 tinding line-line vintages Since in star connection Vab Vine = 03 Vphoge va 4 line voltage Leads Phase voltage 38/ by 30° 1